# Learning Quantum Processes with Quantum Statistical Queries

### v1

@@ Abstract
Learning complex quantum processes is a central challenge in many areas of quantum computing and quantum machine learning, with applications in quantum benchmarking, cryptanalysis, and variational quantum algorithms. This paper introduces the first learning framework for studying quantum process learning within the Quantum Statistical Query (QSQ) model, providing the first formal definition of statistical queries to quantum processes (QPSQs). The framework allows us to propose an efficient QPSQ learner for arbitrary quantum processes accompanied by a provable performance guarantee. We also provide numerical simulations to demonstrate the efficacy of this algorithm. The practical relevance of this framework is exemplified through application in cryptanalysis, highlighting vulnerabilities of Classical-Readout Quantum Physical Unclonable Functions (CR-QPUFs), addressing an important open question in the field of quantum hardware security. This work marks a significant step towards understanding the learnability of quantum processes and shedding light on their security implications.
@@ 1 Introduction
Quantum learning theory constitutes the theoretical framework to study the advantages and limitations of quantum machine learning in the broad sense of it. The main goal is to quantify learning problems as well as design and analyse algorithms to solve them. In this context of computational learning, the learner is a classical, quantum or hybrid algorithm, trying to learn an object (functions, quantum states, quantum processes, distributions, etc.) from a concept class, by having access to some examples of it. Among the different learning models in the classical learning theory literature, some have been extended to quantum and studied under the umbrella of quantum learning theory such as Quantum PAC learning [18, 8], Quantum Statistical Query model [9, 10] and Quantum agnostic learning model [8]. We refer to the following survey for an overview of the results [7]. However more broadly, this framework, in essence, encompasses a wide array of quantum learning results, ranging from learning functions encoded within quantum states [18, 11, 38], quantum state tomography [14, 66, 86], shadow tomography [48, 1] and learning diverse classes of probability distributions [24, 11, 62], to process tomography [61]. While most of the efforts in quantum learning theory have been focused on quantum states (both as examples and target objects), here in this work we shift the focus to quantum processes.
@@
Establishing our framework and definitions for learning quantum processes within the quantum statistical query model enables us to develop efficient learning algorithms for learning general quantum processes under certain distribution of states and conditions on observable. Here, our statistical query learner comes equipped with provable performance guarantees. Furthermore, we explore the practical applications of our model and learning algorithm, venturing into the domain of cryptanalysis. In this domain, our results shed light on the security vulnerabilities of a class of quantum physical unclonable functions, filling a significant gap in the study of these primitives and the protocols employing them.
@@ 1.1 Overview of the main results
  • Framework for learning quantum processes in Quantum Statistical Query model (QPSQ): We provide a formal definition that captures what it means to learn a quantum process using statistical queries. Specifically, we define the quantum statistical query oracle QPStat for a quantum process $\mathcal{E}$ that produces a statistical estimate of the expectation value of an observable $O$, after the application of the quantum process on an input state $\rho$. We show how our oracle generalises the previously defined QStat oracles for states.
@@
  • Efficient QPSQ learner for learning arbitrary quantum processes: We provide an algorithm that can efficiently learn arbitrary processes (under certain conditions on the observable and distribution) in the QPSQ model and hence we provide the first QPSQ learner within our framework. Our algorithm is inspired by the process-learning algorithm of Huang et al. in [46], which itself uses ideas from classical shadow tomography [48]. However, our algorithm, defined within the QPSQ model is applicable more generally and easily to a wider range of scenarios (as shown later in the application section) and provides a more physically-motivated version of the algorithm using similar classical ML methods. We also the query complexity of the algorithm as its provable performance guarantee.
@@
  • Numerical simulations of our learning algorithm: We demonstrate the performance of our proposed algorithm through numerical simulations. We also provide methods for generating valid and simple synthetic statistical queries for our simulations.
@@
  • Application of the QPSQ framework and learner in cryptography: We explore the applications of our model and learning algorithm in cryptanalysis. In this domain, we focus on a primitive from quantum hardware security, namely Classical-Readout of Quantum PUF or CR-QPUF. What makes this primitive a perfect case study for our model is the fact that its main security property relies on the assumption of the inherent difficulty of learning its quantum processes. This assumption has been shown to be violated for a very limited subclass of them with simple underlying unitary circuits using heuristic classical machine learning attacks [69]. However, by employing our efficient learners, we give compelling proof that CR-QPUF with even highly complex underlying quantum circuits with an exponential depth can also be learned in the QPSQ model. Notably, our results show the vulnerability of a large family of protocols relying on CR-QPUFs.
1.2 Related works
@@
The quantum statistical query has been first introduced in [9] where they also show that parity functions, $k$-juntas and polynomial-sized DNF formulas are efficiently learnable in the QSQ model, in contrast to the classical setting where these problems are provably hard. They also introduce the notion of private PAC learning, related to differential privacy, and show its relationship to the QSQ model. The relation between quantum statistical query and differential privacy has been further investigated in [5], studying in particular local differential privacy and QSQ. Recently [10] has dived further into the QSQ model and the role of entanglement in learning in such models and providing QSQ lower bounds on learning quantum states. The paper also compares QSQ to noisy quantum PAC learning, establishing an exponential separation between the two models.
@@
An efficient algorithm for learning the prediction of quantum processes has been introduced in [46] which uses classical shadow [48] as a subroutine. The efficiency guarantee of the algorithm is given for any quantum process under a distribution of input states that is invariant under local Clifford transformation, and with an observable that satisfies certain efficiency and locality conditions. Other approaches for learning unknown quantum processes have been suggested, such as [33], which focuses on learning quantum processes without input control, or [20], which provides methods based on the Pauli transfer matrix for learning quantum processes and Hamiltonians, emphasising the importance of quantum memory.
@@ 1.3 Organisation of the paper
We provide all the necessary background and notations that we use in the paper in Section 11. In Section 3 we introduce our quantum process learning with quantum statistical query (QPSQ) model and we provide the formal definitions and relations to other definitions. In Section 4 we provide our learning algorithm for learning arbitrary processes. We provide our numerical simulations in Section 5 and finally, the applications are discussed in Section 6.
@@ 2.1 Quantum Information
$\mathcal{U}_{N}\mathrel{\mathop{\ordinarycolon}}=\left\{U\in\mathcal{M}_{N,N}\mathrel{\mathop{\ordinarycolon}}UU^{\dagger}=U^{\dagger}U=I\right\}.$ (1)
@@
The set $\mathcal{P}_{1}=\{I,X,Y,Z\}$ forms an orthonormal basis for $\mathcal{M}_{2,2}$ with respect to the Hilbert-Schmidt inner product. The tensor products of Pauli operators and the identity, i.e. the operators of the form $P\in\{I,X,Y,Z\}^{\otimes n}\mathrel{\mathop{\ordinarycolon}}=\mathcal{P}_{n}$, are usually referred as stabilizer operators or Pauli strings over $n$ qubits.
@@
The Haar measure on the unitary group $U(N)$ is the unique probability measure $\mu_{N}$ that is both left and right invariant over the set $\mathcal{U}_{N}$, i.e., for all integrable functions $f$ and for all $V\in\mathcal{U}_{N}$, we have:
@@
Note that the right invariance of the Haar measure implies that the definition of $\mathbb{E}_{\ket{\psi}\sim\mu_{N}}\left[\ket{\psi}\bra{\psi}^{\otimes k}\right]$ does not depend on the choice of |$\ket{\phi}$.
@@
A map $\mathcal{E}$ from one quantum state to another is said to be completely positive if for any positive operator $A,\mathcal{E}(A)$ is also a positive operator. $\mathcal{E}$ is said to be trace-preserving if for any input density operator $\rho,tr(\mathcal{E}(\rho))=1=tr(\rho)$. A quantum process $\mathcal{E}$ is defined as a Completely Positive Trace-Preserving (CPTP) map from one quantum state to another. For a unitary $U$, the associated map is:
:$\mathcal{E}\mathrel{\mathop{\ordinarycolon}}\rho\rightarrow U\rho U^{\dagger}$ (7)
@@ 2.2 Classical Shadow Tomography
Definition 8 (Classical shadow of a quantum state).
@@
Definition 9 (Classical Shadow Approximation of a quantum state).
@@
Definition 10 (Classical shadow of a quantum process).
@@ 2.3 Computational Learning Theory
Computational learning theory studies what it means to learn a function. One of the most successful formal learning frameworks is undoubtedly the model of Probably Approximately Correct (PAC) learning, which was introduced in [82]. In this model, we consider a class of target Boolean functions $\mathcal{C}\subseteq\{f|f\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}\}$, usually called the concept class. For an arbitrary concept $c\in\mathcal{C}$, a PAC learner receives samples of the form $\{x,c(x)\}$, where, in general, $x$ is sampled from an unknown probability distribution $D\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow[0,1]$. In the setting of noisy PAC learning, the bit $c(x)$ of each sample may independently be incorrect with some probability. The learner aims to output, with high probability, a hypothesis function $h$ with low error on average over the distribution $D$.
@@
Quantum generalizations of both PAC and SQ learning have already been introduced and studied widely. The Quantum PAC (QPAC) model was introduced in [18], where the learner has access to a quantum computer and receives quantum example states as input. The quantum example state for a concept $c$ over $n$ input bits with the target distribution $D$ is the $n+1$-qubit state $|\psi_{c}\rangle=\sum_{x}\sqrt{D(x)}|x,c(x)\rangle$. It has been shown in [8] that the sample complexity of quantum and classical PAC learning is the same. However, over a fixed uniform distribution, learning with quantum queries can provide exponential advantage over the classical learner [11, 38].
@@
Let $\mathcal{C}\subseteq\{c\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}\}$ be a concept class and $D\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}$ be a distribution. A quantum statistical query oracle $\mathsf{Qstat}(O,\tau)$ for some $c^{*}\in\mathcal{C}$ receives as inputs $O,\tau$, where $\tau\geq 0$ and $O\in(\mathbb{C}^{2})^{\otimes n+1}\times(\mathbb{C}^{2})^{\otimes n+1},||O||\leq 1$, and returns a number $\alpha$ satisfying
@@
Note that in this QSQ model, while the learner can obtain an estimate of any measurement on the quantum examples, it is restricted only to classical computation. Interestingly, several concept classes such as parities, juntas, and DNF formulae are efficiently learnable in the QSQ model, whereas the classical statistical query model necessitates an exponentially larger number of samples. Additionally, the authors of [10] have established an exponential gap between QSQ learning and learning with quantum examples in the presence of classification noise.

### v3

@@ Abstract
Learning complex quantum processes is a central challenge in many areas of quantum computing and quantum machine learning, with applications in quantum benchmarking, cryptanalysis, and variational quantum algorithms. This paper introduces the first learning framework for studying quantum process learning within the Quantum Statistical Query (QSQ) model, providing the first formal definition of statistical queries to quantum processes (QPSQs). The framework allows us to propose an efficient QPSQ learner for arbitrary quantum processes accompanied by a provable performance guarantee. We also provide numerical simulations to demonstrate the efficacy of this algorithm. In our new framework, we prove exponential query complexity lower bounds for learning unitary 2-designs, and a doubly exponential lower bound for learning haar-random unitaries. The practical relevance of this framework is exemplified through application in cryptography, highlighting vulnerabilities of a large class of Classical-Readout Quantum Physical Unclonable Functions (CR-QPUFs), addressing an important open question in the field of quantum hardware security. This work marks a significant step towards understanding the learnability of quantum processes and shedding light on their security implications.
@@ 1 Introduction
Quantum learning theory constitutes the theoretical framework to study the advantages and limitations of quantum machine learning in the broad sense of it. The main goal is to quantify learning problems as well as design and analyse algorithms to solve them. In this context of computational learning, the learner is a classical, quantum or hybrid algorithm, trying to learn an object (functions, quantum states, quantum processes, distributions, etc.) from a concept class, by having access to some examples of it. Among the different learning models in the classical learning theory literature, some have been extended to quantum and studied under the umbrella of quantum learning theory such as Quantum PAC learning [22, 23], Quantum Statistical Query learning [24, 25] and Quantum agnostic learning [23]. We refer to the following survey for an overview of the results [26]. However more broadly, this framework, in essence, encompasses a wide array of quantum learning results, ranging from learning functions encoded within quantum states [22, 27, 28], quantum state tomography [29, 30, 31], shadow tomography [32, 33] and learning diverse classes of probability distributions [34, 27, 35], to process tomography [36]. While most of the efforts in quantum learning theory have been focused on quantum states (both as examples and target objects), here in this work we shift the focus to quantum processes.
@@
Establishing our framework and definitions for learning quantum processes within the quantum statistical query model enables us to develop efficient learning algorithms for learning general quantum processes under a certain distribution of states and conditions on observable. Here, our statistical query learner comes equipped with provable performance guarantees. Using our new framework, we are also able to provide query complexity lower bounds on the hardness of learning certain classes of unitaries. There are several noteworthy points regarding our learning model. In comparison with conventional quantum process tomography [36], our access model is significantly weaker. In particular, we restrict the learner to measurement statistics on a single copy of the output state, and the learner cannot obtain the results of any entangled measurements with ancillary qubits. Still, studying various learning tasks with this access model is possible. In this work, we study the tasks of learning to predict expectation values of observables on the outcome of an unknown process, as well as the task of process tomography, i.e. approximating the process within bounded diamond distance. Further, we argue that the quantum statistical query access model to a quantum process is a strong formal model to describe learning in practical settings. In many practical scenarios, due to the limitations of quantum memories, only the statistical results of quantum measurements are relevant and accessible in the near term. This aspect is precisely the focus of our model. Furthermore, our model gains importance when considering scenarios in which direct access to the process is not provided. This becomes especially crucial for cryptanalysis purposes, where certain attack models may not grant direct access to the process, but statistical data can still be accessed by adversaries. Our model allows for the study of statistical data, as demonstrated in our application results. Furthermore, we explore the practical applications of our model and learning algorithm and lower bounds, by venturing into the domain of cryptography. In this domain, our results shed light on identifying the conditions under which a class of quantum physical unclonable functions is vulnerable, significantly narrowing the gap in the study of these primitives and the protocols employing them.
@@ 1.1 Overview of the main results
  • Framework for learning quantum processes in Quantum Statistical Query model (QPSQ): We provide a formal definition that captures what it means to learn a quantum process using statistical queries. Specifically, we define the quantum statistical query oracle QPStat for a quantum process $\mathcal{E}$ that produces a statistical estimate of the expectation value of an observable $O$, after the application of the quantum process on an input state $\rho$. We show how our oracle generalises the previously defined QStat oracles for states. We show that in many situations, efficient quantum statistical query algorithms for learning boolean functions admit equivalent, efficient algorithms in our model.
@@
  • Efficient QPSQ learner for learning arbitrary quantum processes: We provide an algorithm that can efficiently learn arbitrary processes (under certain conditions on the observable and states) in the QPSQ model. Our algorithm is inspired by the process-learning algorithm of Huang et al. in [53], which itself uses ideas from classical shadow tomography [32]. We show that even within our weaker access model, we can obtain similar provable performance guarantees with an efficient query complexity, accessing the process only through QPSQs and classical post-processing.
@@
  • Numerical simulations of the learning algorithm: We demonstrate the performance of our proposed algorithm through numerical simulations. We also provide methods for generating valid and simple synthetic statistical queries for our simulations.
  • Hardness of learning classes of unitaries: We provide an exponential lower bound on the query complexity for learning exact and approximate unitary $2$-designs from QPStat queries. We also show a nearly doubly exponential lower bound on the hardness of learning unitaries over the Haar-measure from QPStat queries.
@@
  • Application of the QPSQ framework and learner in cryptography: We explore the applications of our results in cryptography. In this domain, we focus on a primitive from quantum hardware security, namely Classical-Readout of Quantum PUF or CR-QPUF. What makes this primitive a perfect case study for our model is the fact that its main security property relies on the assumption of the inherent difficulty of learning an underlying quantum process. For a very limited subclass of CR-QPUFs, with very simple and uncorrelated underlying unitary circuits, it has been shown that such a learning task is feasible using heuristic classical machine learning attacks [52]. However, due to the lack of any theoretical results for this class, it has been an important open question whether CR-QPUFs with more complicated or deep underlying circuits can provide security. We answer this open question in our work with powerful negative results. Our efficient learning algorithm, which is guaranteed to succeed under certain assumptions, demonstrates the conditions under which such a protocol cannot be secure, showing the vulnerability of a broad family of protocols relying on CR-QPUFs (including many foreseeable practical instances). We are thus able to show that most efficiently implementable CR-QPUFs cannot achieve the desired level of security.
1.2 Related work
@@
The quantum statistical query model was first introduced in [24] where they also show that parity functions, $k$-juntas and polynomial-sized DNF formulas are efficiently learnable in the QSQ model, in contrast to the classical setting where these problems are provably hard. They also introduce the notion of private PAC learning, related to differential privacy, and show its relationship to the QSQ model. The relation between quantum statistical query and differential privacy has been further investigated in [58], studying in particular local differential privacy and QSQ. Recently [25] has dived further into the QSQ model and the role of entanglement in learning in such models and providing QSQ lower bounds on learning quantum states. The paper also compares QSQ to noisy quantum PAC learning, establishing an exponential separation between the two models. An oracle allowing one to make statistical queries to the Choi-state of a process was presented in [59]. The paper then dives into learning algorithms for quantum boolean functions and quantum k-juntas, while also adapting the lower bounds of  [25] to novel problems in this oracle model.
Recent papers that establish lower bounds in quantum statistical query learning, such as [25, 60, 61], derive results by adapting the classical techniques presented by Feldman in [62]. In [60], the authors showed lower bounds for learning the output Born-distributions of random circuits at varying depth regimes. In concurrent work, Nietner in [61] provided a unifying framework for a large class of learning models (including many statistical query oracles), presented lower bounds within this framework, and applied them to the task of learning output states of quantum circuits. Compared to learning within a low diamond distance, the tasks of learning output distributions and output states are easier. However, the access model we consider is also stronger than their statistical query models. As such, our bounds cannot be directly compared to those of [60, 61]. The notion of a ‘multi-copy’ QSQ oracle was also introduced in [61], where one can make statistical queries to many copies of the same state.
@@
An efficient algorithm for learning the prediction of quantum processes using classical shadows [32] has been introduced in [53]. The efficiency guarantee of the algorithm is given for any quantum process under a distribution of input states that is invariant under local Clifford transformations, and with an observable that satisfies certain efficiency and locality conditions. Other approaches for learning unknown quantum processes have been suggested, such as [55], which focuses on learning quantum processes without input control, or [63], which provides methods based on the Pauli transfer matrix for learning quantum processes and Hamiltonians, emphasising the importance of quantum memory.
@@ 1.3 Organisation of the paper
We provide all the necessary background and notations that we use in the paper in Section 14. In Section 3, we formally introduce our model for learning quantum processes with quantum statistical queries (QPSQ) and show its relation to other models. In Section 4, we provide our learning algorithm for learning arbitrary processes, along with numerical simulations in Section 5. In Section 6, we present some query complexity lower bounds in our model. Finally, the applications for cryptography are discussed in Section 7.
@@ 2.1 Quantum Information
$\mathcal{U}_{N}:=\left\{U\in\mathcal{M}_{N,N}:UU^{\dagger}=U^{\dagger}U=I\right\}.$ (1)
@@
The set $\mathcal{P}_{1}=\{I,X,Y,Z\}$ forms an orthonormal basis for $\mathcal{M}_{2,2}$ with respect to the Hilbert-Schmidt inner product. The tensor products of Pauli operators and the identity, i.e. the operators of the form $P\in\{I,X,Y,Z\}^{\otimes n}:=\mathcal{P}_{n}$, are usually referred as stabilizer operators or Pauli strings over $n$ qubits.
@@
The Haar measure on the unitary group $U(N)$ is the unique probability measure $\mu_{H}$ that is both left and right invariant over the set $\mathcal{U}_{N}$, i.e., for all integrable functions $f$ and for all $V\in\mathcal{U}_{N}$, we have:
@@
Note that the right invariance of the Haar measure implies that the definition of $\mathbf{E}_{\ket{\psi}\sim\mu_{S}}\left[\ket{\psi}\bra{\psi}^{\otimes k}\right]$ does not depend on the choice of |$\ket{\phi}$.
@@
A map $\mathcal{E}\in\mathcal{M}_{N,N}$ from one quantum state to another is said to be completely positive if for any positive operator $A\in\mathcal{M}_{N^{2},N^{2}},(\mathcal{E}\otimes I)(A)$ is also a positive operator. $\mathcal{E}$ is said to be trace-preserving if for any input density operator $\rho,Tr(\mathcal{E}(\rho))=1=Tr(\rho)$. A quantum process $\mathcal{E}$ is defined as a Completely Positive Trace-Preserving (CPTP) map from one quantum state to another. For a unitary $U$, the associated map is:
$\mathcal{E}:\rho\rightarrow U\rho U^{\dagger}$ (7)
Definition 6 (Maximally Depolarizing Channel).
The maximally depolarizing channel $\Phi_{dep}$ acting on states in $S_{N}$ is defined as follows:
$\Phi_{dep}(\rho)=\frac{Tr(\rho)}{N}I$ (8)
Definition 7 (Diamond norm and diamond distance).
[…]
$\norm{\mathcal{E}}_{\diamond}=\underset{\rho\in\mathcal{S}_{N^{2}}}{\mathrm{max}}\norm{(\mathcal{E}\otimes I)(\rho)}_{1}$ (9)
We then define the diamond distance $d_{\diamond}$ :
$d_{\diamond}(\mathcal{E}_{1},\mathcal{E}_{2})=\frac{1}{2}\norm{\mathcal{E}_{1}-\mathcal{E}_{2}}_{\diamond}$ (10)
Definition 8 (Choi-Jamiolkowski Isomorphism).
For any linear map $\Phi:\mathcal{M}_{N_{1},N_{1}}\rightarrow\mathcal{M}_{N_{2},N_{2}}$, we define the associated CJ-state $J(\Phi)$:
$J(\Phi):=\overset{N_{1}}{\underset{i,j=1}{\sum}}|i\rangle\langle j|\otimes\Phi(|i\rangle\langle j|)\in\mathcal{M}_{N_{1},N_{1}}\otimes\mathcal{M}_{N_{2},N_{2}}$ (11)
@@ 2.2 Classical Shadow Tomography
Definition 11 (Classical shadow of a quantum state [32]).
@@
Definition 12 (Classical Shadow Approximation of a quantum state [32]).
@@
Definition 13 (Classical shadow of a quantum process [53]).
@@ 2.3 Computational Learning Theory
Computational learning theory studies what it means to learn a function. One of the most successful formal learning frameworks is undoubtedly the model of Probably Approximately Correct (PAC) learning, which was introduced in [67]. In this model, we consider a class of target Boolean functions $\mathcal{C}\subseteq\{f|f:\{0,1\}^{n}\rightarrow\{0,1\}\}$, usually called the concept class. For an arbitrary concept $c\in\mathcal{C}$, a PAC learner receives samples of the form $\{x,c(x)\}$, where, in general, $x$ is sampled from an unknown probability distribution $D:\{0,1\}^{n}\rightarrow[0,1]$. In the setting of noisy PAC learning, the bit $c(x)$ of each sample may independently be incorrect with some probability. The learner aims to output, with high probability, a hypothesis function $h$ with low error on average over the distribution $D$.
@@
Quantum generalizations of both PAC and SQ learning have already been introduced and studied widely. The Quantum PAC (QPAC) model was introduced in [22], where the learner has access to a quantum computer and receives quantum example states as input. The quantum example state for a concept $c$ over $n$ input bits with the target distribution $D$ is the $n+1$-qubit state $|\psi_{c}\rangle=\sum_{x}\sqrt{D(x)}|x,c(x)\rangle$. It has been shown in [23] that the sample complexity of quantum and classical PAC learning is the same. However, over a fixed uniform distribution, learning with quantum queries can provide exponential advantage over the classical learner [27, 28]. Efficient learners from quantum queries under product distributions have also been shown [69, 70].
@@
Let $\mathcal{C}\subseteq\{c:\{0,1\}^{n}\rightarrow\{0,1\}\}$ be a concept class and $D:\{0,1\}^{n}\rightarrow[0,1]$ be a distribution over $n$-bit strings. A quantum statistical query oracle $\mathsf{QStat}(O,\tau)$ for some $c^{*}\in\mathcal{C}$ receives as inputs $O,\tau$, where $\tau\geq 0$ and $O\in(\mathbb{C}^{2})^{\otimes n+1}\times(\mathbb{C}^{2})^{\otimes n+1},||O||\leq 1$, and returns a number $\alpha$ satisfying
@@
Note that in the QSQ model of [24], while the learner can obtain an estimate of any measurement on the quantum examples, it is restricted only to classical computation. Interestingly, several concept classes such as parities, juntas, and DNF formulae are efficiently learnable in the QSQ model, whereas the classical statistical query model necessitates an exponentially larger number of queries. Additionally, the authors of [25] have established an exponential separation between QSQ learning and learning with quantum examples in the presence of classification noise.
One can also define quantum statistical queries for the task of learning states. In this case, the quantum example in definition 14 would be replaced by the unknown quantum state to be learned.